# {tocentry}

### v1

@@ 1 Introduction
Quantum mechanics (QM) is the underlying physical theory of chemical phenomena. However, applying a fully QM treatment to large-scale chemical problems is both unnecessary and computationally prohibitive. Pioneered by Warshel, Levitt, and Karplus1, 2, the hybrid Quantum Mechanics/Molecular Mechanics (QM/MM) approach addresses this by limiting the QM treatment to critical regions of interest, while modeling the surroundings with more approximate empirical force fields3, 4, 5, 6, 7, 8, 9, 10.
In this work, we describe and apply a new QM/MM implementation. Given a QM/MM system partitioning, the factors that define the QM/MM treatment are (i) the choice of method to use in the QM calculation, (ii) the treatment of the electrostatic interaction between the QM and MM regions, (iii) the handling of the QM/MM boundary, and (iv) the quality of the MM force-field. We present a GPU-based QM/MM implementation with the following characteristics (i) fast algorithms for accurate QM methods, such as hybrid density functionals, in the QM region, (ii) a multipole representation and self-consistent treatment of the periodic quantum electrostatic potential (denoted here the QM/MM-Multipole method), (iii) a pseudo-bond treatment of the QM/MM boundary, and an open-source implementation within PySCF taking advantage of the GPU accelerated GPU4PySCF module, thus (iv) enabling straightforward integration with standard open-source force-fields.
We carefully benchmark the convergence of the QM/MM-Multipole method (and its resulting effect on the QM self-consistent field stability and energy conservation) in a QM/MM water model and a model for microtubule-mediated GTP hydrolysis. Finally, we investigate the utility of our implementation in enzyme reactions in the context of the catalytic kinetics of chorismate mutase, that has recently been studied at the QM/MM level using a pure density functional description of the QM region 11. In our simulations, we are able to employ a more accurate QM method obtained by reparametrizing a hybrid density functional to local coupled cluster energetics. Our resulting reaction rate, extracted by enhanced sampling of the QM/MM potential energy surface, improves the agreement with the experimental estimate, and overall our study highlights the sensitivity of the rate to many aspects of the QM/MM treatment.
@@ 2.1 Energy partitioning
Figure 1: An illustration of the partitioning of the QM-MM electrostatics into near-field (region I) and far-field (region II) components. Each rectangle represents one unit cell. Region I is shown as a sphere centered on the QM subsystem, and region II contains the remaining infinite lattice.
@@ 2.2 Electrostatic treatment: background
In condensed phase systems, a critical component of the functional in Eq. 1 is the electrostatic coupling, in particular the contributions that arise from long-range periodic electrostatics within each of the three terms. It has long been recognized that serious artifacts arise in molecular simulations from ignoring the long-range tail of the electrostatics12.
@@
One category of methods represents the electrostatic potential of the periodic MM charge distribution on a grid constructed within the reference cell. The MM electrostatic potential can then be integrated with the QM charge density in the reference cell13, 14. Various kinds of grids and approximations to accelerate this computation have been explored 13, 15, 16, 17, 18, 19.
An alternative approach, named the QM/MM-Ewald method and originally proposed in the context of semi-empirical QM treatments20, 21, 22, 23, 24, 25, 26, 27, 28, approximates the QM electron distribution by QM-atom-centered point charges. The long-range QM/MM electrostatics, i.e. the interaction between the reference cell QM density and periodic MM charges outside of the reference cell, can then be handled using standard point-charge Ewald lattice sum methods, while the interactions between the QM density and nearby MM charges within the reference cell are computed exactly.
The QM/MM-Ewald method has the benefit of simplicity and avoids the costly evaluation of the long-range electrostatic potential on a large number of grid points, but it assumes that the QM charge distribution is well approximated by point charges. This is a good approximation when there is a large separation between the reference cell QM charges and the MM charges in other cells from a large simulation box. However, artifacts have been reported when applying this together with a DFT treatment of the QM region. In particular, difficulties were observed in converging the self-consistent field (SCF) iterations when using Mulliken QM charges and a diffuse basis24, 29, and a stable SCF required refitting the QM charges to the QM electrostatic potential during the SCF cycles24, 27.
@@ 2.3 Multipole Approximation to Electrostatics
Our approach to the long-range electrostatics starts from the simple and efficient QM/MM-Ewald method but extends the QM electronic description to a multipole representation. As we show later in the numerical results, this yields favourable convergence and SCF stability properties without the need to refit the QM charges during the SCF cycles.24, 27
@@
For the QM density contribution of a basis pair $\mu$ and $\nu$ centered on different atoms, the expansion is performed at both centers and then arithmetically averaged. The Taylor series in Eq. 3 has been truncated at second order. The leading error in such an approximation is the missing third-order term that scales asymptotically as $1/|\mathbf{R}^{\mathrm{QM}}_{\mu}-\mathbf{R}^{\mathrm{MM}}_{i}|^{4}$. This is the first term in the series that decays sufficiently fast to ensure its infinite lattice sum (over the periodic MM images) absolutely converges in real space30. This allows for control over the truncation error to achieve the desired accuracy by appropriately selecting Region I. We discuss how this can be accomplished in practice later.
@@
where $Q$, $\mathbf{\mu}$ and $\mathbf{\theta}$ are Mulliken atomic charges, dipoles and quadrupoles31, and the $T_{ji\cdots}$’s are the Taylor expansion coefficients, referred to as the multipole interaction tensors. A similar multipole expansion approximation describes the interactions between the unit-cell QM and periodic QM images, yielding:
@@
We ignore the electronic exchange and correlation effects between QM images, due to their short-range nature. To manage the infinite lattice sum over region II, we utilize the identity $\sum_{i\in\mathrm{II}}=\sum_{i}-\sum_{i\in\mathrm{I}}$, converting the restricted sum over region II into an unrestricted sum over the whole lattice minus a finite real-space sum over region I. We treat the whole lattice sum by the Ewald method for multipoles32 (see also Appendix), and we re-write the QM/MM energy of the periodic system as
@@
where $E^{\mathrm{iso}}_{\mathrm{QM}}$ indicates the QM energy of an isolated QM subsystem i.e. without the periodic QM-QM electrostatic interactions. In Eq. 2.3, the first two terms account for all the QM and QM-MM energies in region I. The third and fourth terms sum to $E_{\mathrm{QM}}^{\mathrm{II}}$, and the fifth and sixth terms sum to $E_{\mathrm{QM-MM}}^{\mathrm{II}}$. The Ewald interaction tensors are denoted as $\psi_{ji\cdots}$, and are given in the Appendix. (When we write a summation range over the unit cell, we are assuming the minimum image convention).
In addition to the MM point charge model, it is useful to consider Gaussian smeared MM charges. Using such Gaussian charges reduces the so-called “electron spill-out” effect20, 17, 7, where density that is very close to a point MM charge is overpolarized. In this context, $E_{\mathrm{QM-MM}}^{\mathrm{I}}$ is calculated as:
@@
where $\hat{T}_{ji\cdots}$ represents the multipole interaction tensors between point charges and Gaussian distributed charges (detailed in the Appendix). Since this correction decays exponentially fast with the charge separation, the infinite lattice sum over MM charges in the above expression reduces to a finite sum in real space, and the truncation can be made based on each MM charge exponent $\kappa_{i}$.
@@ 2.4 Electronic Mean-Field Theory
The nuclear gradient required to perform MD is obtained by differentiating the QM/MM energy with respect to the atom positions. The nuclear gradient of the interaction energy in region I ($E_{\mathrm{QM}}^{\mathrm{iso}}+E_{\mathrm{QM-MM}}^{\mathrm{I}}$) is well-documented in the literature for non-periodic QM/MM systems (see e.g., Ref. 3). The $E^{\mathrm{II}}$ gradient contains both the response of the multipole interaction tensors $T_{ji\cdots}$, $\psi_{ji\cdots}$ and $\hat{T}_{ji\cdots}$, and of the multipole moments $Q$, $\mu$ and $\theta$. The interaction tensors, as purely geometric functions, are independent of the multipole moments, and it is straightforward to differentiate them with respect to the atom positions (see Appendix for detailed expressions). The multipole response formally depends on both a Pulay term (response due to the basis center movement) and the 1-RDM response. However, the 1-RDM response is not explicitly needed if the wavefunction is at a variational stationary point, i.e. as long as the self-consistent HF/KS equations are converged. Given these considerations, the $E^{\mathrm{II}}$ gradient with respect to a QM atom $j$ (denoted $\nabla_{j}E^{\mathrm{II}}$) can be written as
@@ 2.6 Implementation
We have implemented our QM/MM-Multipole method within the PySCF package taking advantage of GPU acceleration via the GPU4PySCF module. Given the full system atom positions, the QM atom element types, and MM charges (and optionally radii), PySCF is tasked with computing $E_{\mathrm{QM}}$, $E_{\mathrm{QM-MM}}$, and their nuclear gradients with respect to all atoms. The computationally expensive SCF calculations are offloaded to GPU routines in GPU4PySCF. The QM-MM electrostatic integral within region I (Eq. 7) is computed using the GPU implementation of the three-center-two-electron integrals in GPU4PySCF. In addition, the interaction tensors $T$, $\hat{T}$, and $\psi$, are computed on the GPU through CuPy48 array operations, while the tensor contractions between the interaction tensors and multipoles are also performed by CuPy. In our QM/MM implementation, it is not necessary to employ a particle-mesh variant of Ewald, which uses a real-space grid to interpolate the charges, to improve the computational scaling. This is because by choosing the Ewald real-space cutoff to be roughly the box length, the computational cost of both the real-space and $k$-space sums scales linearly with the number of MM atoms, since we only compute the QM-QM and QM-MM Ewald energies in GPU4PySCF, and not the MM-MM Ewald energy. Aside from the overlap-like one-electron integrals, used for computing multipole moments and their gradients, which are evaluated on the CPU due to their relatively low cost, all QM/MM-related computational tasks are accelerated by the GPU.
We use LAMMPS49 to compute $E_{\mathrm{MM}}$ and its nuclear gradients. This is straightforwardly achieved within LAMMPS input files, by turning off the bonded interactions within the QM subsystem and the electrostatics between the QM and MM atoms. In our current setup, GPU4PySCF and LAMMPS run concurrently in the background and communicate with an MD integrator, i-PI50. At each MD step, i-PI sends the box dimensions and atom positions via internet or Unix sockets to both GPU4PySCF and LAMMPS and collects the computed energy and forces from them to integrate the nuclear dynamics. Such a communication scheme allows for the overlap of the LAMMPS CPU computation with GPU computation by GPU4PySCF. We note that LAMMPS itself can serve as the MD integrator as well if one desires to eliminate a dependency on i-PI. Presently since there is no communication between GPU4PySCF and LAMMPS, any choice of MD code can be easily adopted into this QM/MM communication scheme and requires no modification if it comes with an out-of-the-box i-PI interface, or if not, only small changes are needed to enable socket communication with i-PI.
@@ 3.1 Common Setup
We outline here computational details that the simulations below share in common. In all the simulations, we modeled the MM charges as Gaussian-distributed charges when computing the QM-MM interactions, with the exponent of the Gaussian on an atom given by the square of the inverse of its covalent radius in Ref. 51. The MM partial charges (used in $E_{\text{QM-MM}}$) and the Lennard-Jones parameters between the QM and MM atoms (used in $E_{\text{MM}}$) were kept the same as in the classical force field used in the simulation. The density fitting approximation to the electron repulsion integrals was employed, using the def2-SVP-JKFIT basis as the auxiliary fitting basis for the def2-SVPD basis, and the 6-311G**-RIFIT basis52 for all the other atomic bases. For the pseudo-bond atoms, the fitting basis was generated by PySCF53 using an even-tempered Gaussian expansion. The SCF convergence criteria were chosen to be $10^{-10}$ Hartrees for the energy and $10^{-6}$ Hartrees for the orbital gradient. In the MD simulations, we used the “time-reversible always stable predictor-corrector” method (ASPC)54, 55 to predict the 1-RDM for the current MD step from the four preceding MD steps. The predicted 1-RDM was used as the initial guess for SCF and for dynamically determining region I based on a given error threshold (detailed further below). In all QM/MM $NVT$ simulations, the temperature was controlled by a Langevin thermostat with a damping time constant of 100 fs. A time step of 0.5 fs was used for all QM/MM MD unless stated otherwise.
@@ 3.2 QM Water in MM Water
For this application, we considered three QM-water-in-MM-water systems, consisting of 7 QM water molecules embedded in respectively 1, 221, and 33234 SPC/Fw56 MM water molecules. The initial configuration of the 221 and 33234 MM water systems was generated using Packmol57 by placing 228/33241 water molecules in cubic boxes with 19.002 Å and 100.013 Å side lengths, according to a 0.99403 kg/L water density. The system was energy minimized and equilibrated for 50 ps at 298.15 K at the SPC/Fw MM level. The temperature was controlled by a Nose-Hoover chain and the dynamics were integrated using a 0.5 fs time step. The final equilibrated configuration was used for subsequent QM/MM simulations, where the 7 closest waters to the box center were selected as QM waters. All QM/MM calculations employed the PBE58/def2-SVPD level of theory. The 7-QM-water-in-1-MM-water system was built by extracting the QM water configuration from the equilibrated 228-water structure and placing it at the center of a cubic box with a 27.399 Å side length. An MM water molecule was then positioned along one of the box diagonals with varying distances from the box center.

### v2

@@ 1 Introduction
Quantum mechanics (QM) is the underlying physical theory of chemical phenomena. However, applying a full QM treatment to large-scale chemical problems is both unnecessary and computationally prohibitive. Pioneered by Warshel, Levitt, and Karplus1, 2, the hybrid Quantum Mechanics/Molecular Mechanics (QM/MM) approach addresses this by limiting the QM treatment to critical regions of interest while modeling the surroundings with more approximate empirical force fields3, 4, 5, 6, 7, 8, 9, 10.
In this work, we describe and apply a new QM/MM implementation. Given a QM/MM system partitioning, the factors that define the QM/MM treatment are (i) the choice of method to use in the QM calculation, (ii) the treatment of the electrostatic interaction between the QM and MM regions, (iii) the handling of the QM/MM boundary, and (iv) the quality of the MM force-field. We present a GPU-based QM/MM implementation with the following characteristics (i) fast algorithms for accurate QM methods, such as hybrid density functionals, in the QM region, (ii) a multipole representation and self-consistent treatment of the periodic quantum electrostatic potential (denoted here the QM/MM-Multipole method), (iii) a pseudo-bond treatment of the QM/MM boundary, and an open-source implementation within PySCF taking advantage of the GPU accelerated GPU4PySCF module11, 12, thus (iv) enabling straightforward integration with standard open-source force-fields.
We carefully benchmark the convergence of the QM/MM-Multipole method (and its resulting effect on the QM self-consistent field stability and energy conservation) in a QM/MM water model and a model for microtubule-mediated GTP hydrolysis. Finally, we investigate the utility of our implementation in enzyme reactions in the context of the catalytic kinetics of chorismate mutase, which has recently been studied at the QM/MM level using a pure density functional description of the QM region 13. In our simulations, we are able to employ a more accurate QM method obtained by reparametrizing a hybrid density functional to local coupled cluster energetics. Our resulting reaction rate, extracted by enhanced sampling of the QM/MM potential energy surface, improves the agreement with the experimental estimate, and overall our study highlights the sensitivity of the rate to many aspects of the QM/MM treatment.
@@ 2.1 Energy partitioning
Figure 1: An illustration of the partitioning of the QM-MM electrostatics into near-field (region I) and far-field (region II) components. Each rectangle represents one unit cell. Region I is shown as a sphere centered on the QM subsystem and region II contains the remaining infinite lattice.
@@ 2.2 Electrostatic treatment: background
In condensed phase systems, a critical component of the functional in Eq. 1 is the electrostatic coupling, in particular the contributions that arise from long-range periodic electrostatics within each of the three terms. It has long been recognized that serious artifacts arise in molecular simulations from ignoring the long-range tail of the electrostatics14, 15.
@@
One category of methods represents the electrostatic potential of the periodic MM charge distribution on a grid constructed within the reference cell. The MM electrostatic potential can then be integrated with the QM charge density in the reference cell16, 17. Various kinds of grids and approximations to accelerate this computation have been explored 16, 18, 19, 20, 21, 22.
An alternative approach, named the QM/MM-Ewald method and originally proposed in the context of semi-empirical QM treatments23, 24, 25, 26, 27, 28, 29, 30, 31, approximates the QM electron distribution by QM-atom-centered point charges. The long-range QM/MM electrostatics, i.e. the interaction between the reference cell QM density and periodic MM charges outside of the reference cell, can then be handled using standard point-charge Ewald lattice sum methods, while the interactions between the QM density and nearby MM charges within the reference cell are computed exactly.
The QM/MM-Ewald method has the benefit of simplicity and avoids the costly evaluation of the long-range electrostatic potential on a large number of grid points, but it assumes that the QM charge distribution is well approximated by point charges. This is a good approximation when there is a large separation between the reference cell QM charges and the MM charges in other cells from a large simulation box. However, artifacts have been reported when applying this together with a DFT treatment of the QM region. In particular, difficulties were observed in converging the self-consistent field (SCF) iterations when using Mulliken QM charges and a diffuse basis27, 32, and a stable SCF required refitting the QM charges to the QM electrostatic potential during the SCF cycles27, 30.
@@ 2.3 Multipole Approximation to Electrostatics
Our approach to the long-range electrostatics starts from the simple and efficient QM/MM-Ewald method but extends the QM electronic description to a multipole representation33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45. As we show later in the numerical results, this yields favorable convergence and SCF stability properties without the need to refit the QM charges during the SCF cycles.27, 30
@@
For the QM density contribution of a basis pair $\mu$ and $\nu$ centered on different atoms, the expansion is performed at both centers and then arithmetically averaged. The Taylor series in Eq. 3 has been truncated at the second order. The leading error in such an approximation is the missing third-order term that scales asymptotically as $1/|\mathbf{R}^{\mathrm{QM}}_{\mu}-\mathbf{R}^{\mathrm{MM}}_{i}|^{4}$. This is the first term in the series that decays sufficiently fast to ensure its infinite lattice sum (over the periodic MM images) absolutely converges in real space46. This allows for control over the truncation error to achieve the desired accuracy by appropriately selecting Region I. We discuss how this can be accomplished in practice later.
@@
where $Q$, $\mathbf{\mu}$ and $\mathbf{\theta}$ are Mulliken atomic charges, dipoles and quadrupoles47, and the $T_{ji\cdots}$’s are the Taylor expansion coefficients, referred to as the multipole interaction tensors. A similar multipole expansion approximation describes the interactions between the unit-cell QM and periodic QM images, yielding:
@@
We ignore the electronic exchange and correlation effects between QM images, due to their short-range nature. To manage the infinite lattice sum over region II, we utilize the identity $\sum_{i\in\mathrm{II}}=\sum_{i}-\sum_{i\in\mathrm{I}}$, converting the restricted sum over region II into an unrestricted sum over the whole lattice minus a finite real-space sum over region I. We treat the whole lattice sum by the Ewald method for multipoles48 (see also SI), and we re-write the QM/MM energy of the periodic system as
@@
where $E^{\mathrm{iso}}_{\mathrm{QM}}$ indicates the QM energy of an isolated QM subsystem i.e. without the periodic QM-QM electrostatic interactions. In Eq. 2.3, the first two terms account for all the QM and QM-MM energies in the region I. The third and fourth terms sum to $E_{\mathrm{QM}}^{\mathrm{II}}$, and the fifth and sixth terms sum to $E_{\mathrm{QM-MM}}^{\mathrm{II}}$. The Ewald interaction tensors are denoted as $\psi_{ji\cdots}$ and are given in the SI. (When we write a summation range over the unit cell, we are assuming the minimum image convention).
In addition to the MM point charge model, it is useful to consider Gaussian smeared MM charges. Using such Gaussian charges reduces the so-called “electron spill-out” effect23, 20, 7, where density that is very close to a point MM charge is overpolarized. In this context, $E_{\mathrm{QM-MM}}^{\mathrm{I}}$ is calculated as:
@@
where $\hat{T}_{ji\cdots}$ represents the multipole interaction tensors between point charges and Gaussian distributed charges (detailed in the SI). Since this correction decays exponentially fast with the charge separation, the infinite lattice sum over MM charges in the above expression reduces to a finite sum in real space, and the truncation can be made based on each MM charge exponent $\kappa_{i}$.
@@ 2.4 Electronic Mean-Field Theory
The nuclear gradient required to perform MD is obtained by differentiating the QM/MM energy with respect to the atom positions. The nuclear gradient of the interaction energy in the region I ($E_{\mathrm{QM}}^{\mathrm{iso}}+E_{\mathrm{QM-MM}}^{\mathrm{I}}$) is well-documented in the literature for non-periodic QM/MM systems (see e.g., Ref. 3). The $E^{\mathrm{II}}$ gradient contains both the response of the multipole interaction tensors $T_{ji\cdots}$, $\psi_{ji\cdots}$ and $\hat{T}_{ji\cdots}$, and of the multipole moments $Q$, $\mu$ and $\theta$. The interaction tensors, as purely geometric functions, are independent of the multipole moments, and it is straightforward to differentiate them with respect to the atom positions (see SI for detailed expressions). The multipole response formally depends on both a Pulay term (response due to the basis center movement) and the 1-RDM response. However, the 1-RDM response is not explicitly needed if the wavefunction is at a variational stationary point, i.e. as long as the self-consistent HF/KS equations are converged. Given these considerations, the $E^{\mathrm{II}}$ gradient with respect to a QM atom $j$ (denoted $\nabla_{j}E^{\mathrm{II}}$) can be written as
@@ 2.6 Implementation
We have implemented our QM/MM-Multipole method within the PySCF package taking advantage of GPU acceleration via the GPU4PySCF module. Given the full system atom positions, the QM atom element types, and MM charges (and optionally radii), PySCF is tasked with computing $E_{\mathrm{QM}}$, $E_{\mathrm{QM-MM}}$, and their nuclear gradients with respect to all atoms. The computationally expensive SCF calculations are offloaded to GPU routines in GPU4PySCF. The QM-MM electrostatic integral within region I (Eq. 7) is computed using the GPU implementation of the three-center-two-electron integrals in GPU4PySCF. In addition, the interaction tensors $T$, $\hat{T}$, and $\psi$, are computed on the GPU through CuPy64 array operations, while the tensor contractions between the interaction tensors and multipoles are also performed by CuPy. In our QM/MM implementation, it is not necessary to employ a particle-mesh variant of Ewald, which uses a real-space grid to interpolate the charges, to improve the computational scaling. This is because by choosing the Ewald real-space cutoff to be roughly the box length, the computational cost of both the real-space and $k$-space sums scales linearly with the number of MM atoms since we only compute the QM-QM and QM-MM Ewald energies in GPU4PySCF, and not the MM-MM Ewald energy. Aside from the overlap-like one-electron integrals, used for computing multipole moments and their gradients, which are evaluated on the CPU due to their relatively low cost, all QM/MM-related computational tasks are accelerated by the GPU.
We use LAMMPS65 to compute $E_{\mathrm{MM}}$ and its nuclear gradients. This is straightforwardly achieved within LAMMPS input files, by turning off the bonded interactions within the QM subsystem and the electrostatics between the QM and MM atoms. In our current setup, GPU4PySCF and LAMMPS run concurrently in the background and communicate with an MD integrator, i-PI66. At each MD step, i-PI sends the box dimensions and atom positions via internet or Unix sockets to both GPU4PySCF and LAMMPS and collects the computed energy and forces from them to integrate the nuclear dynamics. Such a communication scheme allows for the overlap of the LAMMPS CPU computation with GPU computation by GPU4PySCF. We note that LAMMPS itself can serve as the MD integrator as well if one desires to eliminate a dependency on i-PI. Presently since there is no communication between GPU4PySCF and LAMMPS, any choice of MD code can be easily adopted into this QM/MM communication scheme and requires no modification if it comes with an out-of-the-box i-PI interface, or if not, only small changes are needed to enable socket communication with i-PI.
@@ 3.1 Common Setup
We outline here computational details that the simulations below share in common. In all the simulations, we modeled the MM charges as Gaussian-distributed charges when computing the QM-MM interactions, with the exponent of the Gaussian on an atom given by the square of the inverse of its covalent radius in Ref. 67. The MM partial charges (used in $E_{\text{QM-MM}}$) and the Lennard-Jones parameters between the QM and MM atoms (used in $E_{\text{MM}}$) were kept the same as in the classical force field used in the simulation. We adopted an Ewald real-space cutoff roughly the size of the box length and estimated the Ewald parameters according to Ref. 68 given a desired $10^{-8}$ Hartree accuracy. The density fitting approximation to the electron repulsion integrals was employed, using the def2-SVP-JKFIT basis as the auxiliary fitting basis for the def2-SVPD basis, and the 6-311G**-RIFIT basis69 for all the other atomic bases. For the pseudo-bond atoms, the fitting basis was generated by PySCF70 using an even-tempered Gaussian expansion. The SCF convergence criteria were chosen to be $10^{-10}$ Hartrees for the energy and $10^{-6}$ Hartrees for the orbital gradient. In the MD simulations, we used the “time-reversible always stable predictor-corrector” method (ASPC)71, 72 to predict the 1-RDM for the current MD step from the four preceding MD steps. The predicted 1-RDM was used as the initial guess for SCF and for dynamically determining region I based on a given error threshold (detailed further below). In all QM/MM $NVT$ simulations, the temperature was controlled by a Langevin thermostat with a damping time constant of 100 fs. A time step of 0.5 fs was used for all QM/MM MD unless stated otherwise.
@@ 3.2 QM Water in MM Water
For this application, we considered three QM-water-in-MM-water systems, consisting of 7 QM water molecules embedded in respectively 1, 221, and 33234 SPC/Fw73 MM water molecules. The initial configuration of the 221 and 33234 MM water systems was generated using Packmol74 by placing 228/33241 water molecules in cubic boxes with 19.002 Å and 100.013 Å side lengths, according to a 0.99403 kg/L water density. The system was energy minimized and equilibrated for 50 ps at 298.15 K at the SPC/Fw MM level. The temperature was controlled by a Nose-Hoover chain and the dynamics were integrated using a 0.5 fs time step. The final equilibrated configuration was used for subsequent QM/MM simulations, where the 7 closest waters to the box center were selected as QM waters. All QM/MM calculations employed the PBE75/def2-SVPD level of theory. The 7-QM-water-in-1-MM-water system was built by extracting the QM water configuration from the equilibrated 228-water structure and placing it at the center of a cubic box with a 27.399 Å side length. An MM water molecule was then positioned along one of the box diagonals with varying distances from the box center.